OpenStudy (anonymous):

289.25 is 89% of what number

OpenStudy (amigatour):

New/original=P/100

OpenStudy (amigatour):

289.25/A=89/100=325

OpenStudy (mathteacher1729):

Another way to write "289.25 is 89% of what number" is to say "89 percent of x is 289.25, find x." Writing that as an equation you get: \[(0.89)(x)=289.25\] Now do you know what to do?

OpenStudy (amigatour):

That's actually better than what I have posted.

OpenStudy (amigatour):

When I learned basic math, I was taught the long ways. The Mathteacher's formula is what I have been using after passing Math101. It's a lot easier.